# Efficient Neural Combinatorial Optimization Solver for the Min-max Heterogeneous Capacitated Vehicle Routing Problem

### v1

@@ V-C Generalization Studies
V-D Ablation Studies
In this subsection, we conduct ablation studies to investigate the contributions of key ECHO design components. The performance are evaluated on MMHCVRP instances with the number of three vehicles and 60 nodes, and the results are presented in Table IV. Specifically, in w/o dual-modality encoder, the node encoder processes only node attributes, ignores edge attributes, and disables the cross‑attention operator used to fuse edge and node features. In w/o PFCA mechanism, the decoder omits historical vehicle information by disabling the PFCA mechanism. In w/o vehicle augment, only node data are augmented following the prior study [44], while vehicle data remain unchanged. As shown in Table IV, the experimental results demonstrate that each component contributes to the SOTA performance of ECHO, particularly the PFCA mechanism, which leverages historically selected vehicles to prevent myopic decision making.
TABLE IV: Ablation Study Results on Different Design Choices
@@ V-D Ablation Studies
w/o dual-modality encoder 1.06
w/o PFCA mechanism 1.16
w/o vehicle augment 1.04

### v2

@@ 5.3 Generalization Studies
5.4 Ablation Studies
In this subsection, we conduct ablation studies to investigate the contributions of key ECHO design components. The performance is evaluated on MMHCVRP instances with the number of three vehicles and 60 nodes, and the results are presented in Table 6. Specifically, in w/o dual-modality encoder, the node encoder processes only node attributes, ignores edge attributes, and disables the cross‑attention operator used to fuse edge and node features. In w/o PFCA mechanism, the decoder omits historical vehicle information by disabling the PFCA mechanism. In w/o vehicle augment, only node data are augmented following the prior study [pomo], while vehicle data remain unchanged. As shown in Table 6, the experimental results demonstrate that each component contributes to the SOTA performance of ECHO.
Table 6: Ablation Study Results on Different Design Choices
@@ 5.4 Ablation Studies
w/o dual-modality encoder 1.82
w/o PFCA mechanism 1.28
w/o vehicle augment 1.20
@@ 6 Conclusion
7 Acknowledgments
This work is supported in part by the Jilin Provincial Department of Science and Technology Project under Grant 20240101369JC.